OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 8y = 8 x + 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (inkyvoyd):

multiply the second equation by 4. You should see an interesting result.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

i mean B

OpenStudy (inkyvoyd):

Are you sure?

OpenStudy (anonymous):

4x + 8y = 8 x + 2y = 2 4 4 x + 8y = 8

OpenStudy (anonymous):

8/4=2 not exactly sure if im right

OpenStudy (inkyvoyd):

you forgot to multiply the x The equation you get is the same as the first one. 4x + 8y = 8 x + 2y = 2 -> 4x+8y=4

OpenStudy (inkyvoyd):

What does this mean?

OpenStudy (anonymous):

oh